# Motion Prompting: Controlling Video Generation with Motion Trajectories

### v1

@@ Abstract
00footnotetext: Work done as an intern, Project lead
@@ 2 Related Work
Recent work explores sparse (and often object-centric) trajectory-based motion control. The approaches vary in their design choices but often require certain complicated engineering techniques for stable training and better convergence. Tora [86], MotionCtrl [73], DragNUWA [81], Image Conductor [37], and MCDiff [8] adopt two-stage (e.g., finetuning with first dense and then sparse trajectories, or training adapters sequentially), specialized losses [37, 43], or multi-stage fine-tuning for multiple modules [71, 9, 59]. MOFA-Video [44] requires separate adapters for different motion types, TrackGo [89] uses custom losses and layers, while other works [81, 73, 44, 37, 86] engineer data filtering pipelines. In contrast, we find that a simpler training recipe yields high quality results. Our model is trained in a single stage, with uniformly sampled dense trajectories, and without any specialized engineering efforts. Yet it handles a wide range of tasks and motions, generalizing to both sparse and dense trajectories during inference.
@@ 3.2 Architecture
for all tracks at each timestep $t$, where multiplying by the visibility $\mathbf{v}[n,t]$ zeros out the embedding if the track is not visible at that location and time. We quantize $x_{t}^{n}$ and $y_{t}^{n}$ to the nearest integer for simplicity. When multiple tracks pass through the same space-time location, we add the embeddings together. The track embeddings $\phi^{n}$ are randomly (regardless of spatial location) assigned sinusoidal positional embeddings from [68]. For completely dense tracks, this representation is equivalent to starting with a dense grid of embeddings and forward warping, similar to [57].
@@ 4 Motion Prompts
Emergent Phenomena. We find that these “interaction” motion prompts can result in straightforward motions, such as turning the head of a parrot in Fig. 3a. But interestingly, we also observe more complex dynamics: e.g., in Fig. 3b, where the tracks toss the hair of the subject, or in Fig. 3d where the sand is swept around. In these examples, we are essentially probing the video prior learned by the model, and by doing so are able to visualize the physics and general world understanding that the model has learned. This suggests two things. First, that we may debug video models with motion prompts and find their limitations, which we discuss further in Sec. 4.6. And second, that as these models improve, we might be able to use these motion prompts to query video generation models as world models that react to user interactions.
@@ 6 Conclusion
Acknowledgements We would like to thank Sarah Rumbley, Roni Paiss, Jeong Joon Park, Liyue Shen, Stella Yu, Alyosha Efros, Daniel Watson, David Fleet, and Bill Freeman for their invaluable feedback and discussions.
@@ A.9 Human Studies
To perform the human studies, we handcraft 30 inputs with diverse image subjects and input motions. Motions consist of a single uninterrupted trajectory. Text prompts are designed to describe the image, but not the desired motion, to factor out the influence of text as much as possible. DragAnything requires masks, which we obtain by running SAM [36] on the first frame with the initial location of the track as the query point. For our method, we turn the trajectory into a grid of tracks as described above. We then feed these inputs to the models and take a single random sample. We follow the same protocol as above for Image Conductor. This results in 30 samples for each model and 90 samples in total.
@@ Appendix C Full Quantitative Results
Appendix D Motion Magnification

### v2

@@ Abstract
00footnotetext: Work done as intern, Project lead
@@ 2 Related Work
Early work proposes video control through sparse motions [21, 2]. More recent work explores similar ideas with more powerful models. The approaches vary in their design choices but often require certain complicated engineering techniques for stable training and better convergence. Tora [89], MotionCtrl [75], DragNUWA [84], Image Conductor [39], and MCDiff [9] adopt two-stage (e.g., finetuning with first dense and then sparse trajectories, or training adapters sequentially), specialized losses [39, 45], architectures [80, 17], or multi-stage fine-tuning for multiple modules [73, 9, 61]. MOFA-Video [46] requires separate adapters for different motion types, TrackGo [92] uses custom losses and layers, while other works [84, 75, 46, 39, 89] engineer data filtering pipelines. In contrast, we find that a simpler training recipe yields high quality results. Our model is trained in a single stage, with uniformly sampled dense trajectories, and without any specialized engineering efforts. Yet it handles a wide range of tasks and motions, generalizing to both sparse and dense trajectories during inference.
@@ 3.2 Architecture
for all tracks at each timestep $t$, where multiplying by the visibility $\mathbf{v}[n,t]$ zeros out the embedding if the track is not visible at that location and time. We quantize $x_{t}^{n}$ and $y_{t}^{n}$ to the nearest integer for simplicity. When multiple tracks pass through the same space-time location, we add the embeddings together. The track embeddings $\phi^{n}$ are randomly drawn from a fixed pool, and act simply as a unique identifier for each track. For completely dense tracks, this representation is equivalent to starting with a dense grid of embeddings and forward warping, similar to [59].
@@ 4 Motion Prompts
Emergent Phenomena. We find that these “interaction” motion prompts can result in straightforward motions, such as turning the head of a parrot in Fig. 3a. But interestingly, we also observe more complex dynamics: e.g., in Fig. 3b, where the tracks toss the hair of the subject, or in Fig. 3d where the sand is swept around. In these examples, we are essentially probing the video prior learned by the model, and by doing so are able to visualize the physics and general world understanding that the model has learned. Furthermore, because our model supports temporally sparse track conditioning, we can effectively do prediction. That is, we can query our model with a motion for a short duration, and then let the model predict the future, allowing us to answer questions such as ”how will the hair behave if I pull on it this way or that?” as in Fig. 3b.
@@ 6 Conclusion
Acknowledgements We would like to thank Sarah Rumbley, Roni Paiss, Jeong Joon Park, Liyue Shen, Stella Yu, Alyosha Efros, Boqing Gong, Daniel Watson, David Fleet, and Bill Freeman for their invaluable feedback and discussions.
@@ A.9 Human Studies
To perform the human studies, we handcraft 30 inputs with diverse image subjects and input motions. Motions consist of a single uninterrupted trajectory. Text prompts are designed to describe the image, but not the desired motion, to factor out the influence of text as much as possible. DragAnything requires masks, which we obtain by running SAM [38] on the first frame with the initial location of the tracks as query points. For our method, we turn the trajectory into a grid of tracks as described above. We then feed these inputs to the models and take a single random sample. We follow the same protocol as above for Image Conductor. This results in 30 samples for each model and 90 samples in total.
@@ Appendix C Full Quantitative Results
Appendix D Human Pose Control
We show results on using our method to control humans through human pose estimated keypoints in Fig. A3. To do this, we first estimate the pose with an off-the-shelf model, and then apply motions to desired keypoints and feed it to our model.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure A3: Pose Conditioning. We estimate human pose, animate it, translate it to tracks, and then feed it to our model. In each row, we show frames from generated videos with input tracks overlaid on top.
Appendix E Motion Magnification